OpenStudy (anonymous):

\[\int\limits_{0}^{1} x^2 (x^2 + 1)^3 dx\] find the antiderivative please

OpenStudy (anonymous):

those are all powers of 3 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\int\limits_{0}^{1}x ^{3}(x ^{3}+1)(x ^{3}+1))x ^{3}+1)dx \rightarrow \int\limits_{0}^{1}(x ^{6}+x ^{3})(x ^{3}+1)(x ^{3}+1) \]

OpenStudy (anonymous):

\[\int\limits_{0}^{1}(x ^{9}+2x ^{6}+x ^{3})(x ^{3}+1)dx \rightarrow \int\limits_{0}^{1}(x ^{12}+3x ^{9}+3x ^{6}+x ^{3})dx\]

OpenStudy (anonymous):

\[\frac{ x ^{13} }{ 13 }+\frac{ 3x ^{10} }{ 10 }+\frac{ 3x ^{7} }{ 7 }+\frac{ x ^{4} }{ 4 } = \frac{ 1 }{ 13 }+\frac{ 3 }{ 10 }+\frac{ 3 }{ 7 }+\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

you see everything? Just tedious.

OpenStudy (anonymous):

And I also didnt feel like using integration by parts.

OpenStudy (amistre64):

expanding into a poly works ... time consuming, but it works ;)

OpenStudy (anonymous):

amistre is there another way to do it besides integration by parts?

OpenStudy (amistre64):

im sure there is, but if they are asking for an "antiderivative", then your setup was what they expected